OpenStudy (mathhelp346):

HELP PLEASE!!!: f(x)=3x-4 g(x)=-x^2 1.(f o g)(-2) 2.(f o f)(-2) f(x)=2x g(x)=x^2+2 h(x)=-4x+3 1.(h o f)o g

OpenStudy (mathhelp346):

for the first part, what does the 2 do? i know how to solve (f o g)(x)

OpenStudy (anonymous):

(f o f)(-2) is basically saying f( f(-2) ), in other words, in 3x-4, replace x with ( 3x - 4 ) so you get f o f(-2) = 3( 3(-2) - 4 ) - 4 then simplify.

OpenStudy (mathhelp346):

Oh ok

OpenStudy (mathhelp346):

and how would i do the second part?

OpenStudy (mathhelp346):

like h(f(x)(g(x))? but then how do i solve it

OpenStudy (anonymous):

(h o f)(x) = h(f(x)) so (h o f)o g = h(f(x)) o g = h( f ( g(x) ) ) f(g(x)) = 2(x^2+2) so h ( f ( g ( x ) ) ) = -4(2(x^2+2) )+3 Is this clear enough?

OpenStudy (mathhelp346):

yes thank you

OpenStudy (anonymous):

Glad I helped.

OpenStudy (mathhelp346):

wait but if f o g would nt it be h(f(x)(g(x))

OpenStudy (mathhelp346):

h o f is h(f(x)) and then there's o g so wouldnt it be h(f(x)(g(x))

OpenStudy (anonymous):

Well, h o f o g = h( f( g(x) ) ) Basically (h o f) o g is same since h and f go first anyway.

OpenStudy (mathhelp346):

ah ok thanks